write in standard form an equation of the line that passes through the given point and has the given slope.use integer coefficients. (-2,-1),m=3
Answer by rapaljer(4671)

y-y1 = m(x-x1)
y-(-1) = 3(x-(-2))
y+1 = 3(x+2)
y+1=3x+6

Subtract y from each side:
1=3x-y +6

Subtract 6 from each side:
1-6=3x-y
-5=3x-y
3x-y=-5

Here is a NEAT trick! Check the answer by substituting the point (-2,-1) to see if it works:
3x-y=-5
3(-2) -(-1) = -5
-6 + 1 = -5
IT CHECKS!!

Dr. Rapalje, Retired
Seminole State College of Florida